OpenStudy (anonymous):

verify: sin^4(x)-cos^4(x)=2sin^2(x)-1

OpenStudy (dumbcow):

\[\cos ^{4} x = \cos ^{2} x * \cos ^{2} x\] \[\cos ^{2} x = 1-\sin ^{2} x\] \[(1-\sin ^{2}x)^{2} = \sin ^{4} x -2\sin ^{2}x +1\]